OpenStudy (anonymous):

b.The time it takes to multiply two matrices of size n is determined by the polynomial expression 2n3 – n2. For example, it takes 45 operations to multiply two 3x3 matrices. Compute the number of operations required to run a program that multiplies two matrices of size 20.

OpenStudy (anonymous):

2*20^3 - 20^2 = 2*8000 - 400 = 15,600 operations

OpenStudy (anonymous):

woot thanks!!

OpenStudy (anonymous):

yw